# AI-Assisted Game Management Decisions: A Fuzzy Logic Approach to Real-Time Soccer Substitutions

### v2

@@ Abstract
In elite soccer, substitution decisions entail significant financial and sporting consequences yet remain heavily reliant on intuition or predictive models that merely mimic historical biases. This paper introduces a Fuzzy Logic-based Decision Support System (DSS) designed for real-time, prescriptive game management. Unlike traditional Machine Learning approaches that encounter a “predictive ceiling” by attempting to replicate human behavior, our system audits performance through an objective, rule-based inference engine. We propose a methodological advancement by reformulating the PlayeRank metric into a Cumulative Mean with Role-Aware Normalization, eliminating the play-time exposure bias inherent in cumulative sum models to enable accurate intra-match comparison. The system integrates this refined metric with physiological proxies (fatigue) and contextual variables (disciplinary risk modulated by tactical role) to calculate a dynamic Substitution Priority ($P_{final}$). Validation via a case study of the 2018 FIFA World Cup match between Brazil and Belgium demonstrates the system’s ecological validity: it not only aligned with expert consensus on executed substitutions (e.g., Gabriel Jesus) but, crucially, identified high-risk scenarios ignored by human decision-makers. Specifically, the model flagged the “Fagner Paradox”—a maximum-priority defensive risk—minutes before a critical yellow card, and detected the “Lukaku Paradox,” where an isolated assist masked a severe drop in participation. These results confirm that Fuzzy Logic offers a transparent, explainable, and superior alternative to “black-box” models for optimizing real-time tactical decisions.
@@ 2 Literature Review
Recent work supports this perspective. SchmidtLilloBustos2024 demonstrate that player influence fluctuates meaningfully over a match due to tactical reconfiguration, opponent pressure, and physical factors, advocating for temporal windowing to capture such dynamics. Following this insight, the present work adopts 5-minute temporal slices, but advances the methodology by replacing the cumulative sum with a cumulative mean across slices. This modification preserves interpretability while enabling direct comparability between players with different playing times, eliminating residual exposure bias and allowing the performance curve to reflect true temporal variability.
A second methodological challenge concerns the interpretation of imprecise or uncertain information in decision-making contexts. The literature highlights fuzzy logic as an effective tool for this purpose. For instance, Huarachi-Macuri et al. LACCEI2023 demonstrate the suitability of fuzzy systems for transforming quantitative descriptors into interpretable linguistic assessments (e.g., Stamina, Agility). Within tactical modeling, Marliere2017 uses fuzzy control systems to arbitrate between tactical states under uncertainty. These contributions motivate the role of fuzzy logic in the present system: integrating performance indicators and contextual factors into a Substitution Priority score in a manner compatible with the inherent uncertainty of in-game decision making.
@@
3 Dataset
@@
4 Data Integration and Pre-processing
@@
4.1 Stage 1: Temporal Performance Metric Generation
@@
$\mathrm{playerank\_acumulativo\_media\_raw}=\frac{1}{t}\sum_{i=1}^{t}\mathrm{playerank\_fatia\_raw}^{(i)}$
@@
4.2 Stage 2: Contextual and Demographic Enrichment
@@
  • momentum_rate: a measure of short-term performance trend, calculated as the difference between the the current($t$) and the previous(t1t-1) 5-minute
    playerank_acumulativo_media_percentil slices.
@@
4.3 Stage 3: Final Cleaning and Role-Aware Normalization
@@
  • removal of out-of-field periods: Periods where it was identified that the player was substituted, or entered as a substitute, were removed to ensure that the metric remained accurate, not applied to goalkeepers
@@ 4.4 Data Validation
5 Exploratory Data Analysis
@@ 5.1 Descriptive Overview
(playerank_acumulativo_media_percentil) 0.501 symmetrical distribution
@@ 5.2 Univariate Distributions
5.3 Role-Based Event Distributions
To further interpret positional behavior, the dataset was aggregated by player position to estimate empirical probabilities of key in-game events (goals, assists, and yellow cards). The empirical distributions align with established tactical expectations:
@@ 5.3 Role-Based Event Distributions
5.4 Variable Definition, Independence and Correlations
5.4.1 Variable Definition and Selection
@@
Dimension Variable (ID) Mathematical Set Description and Relevance
Technical Performance P_cum $0,1]\subset\mathbb{R}$ Cumulative performance percentile per match and position; main indicator of global efficiency.
Fatigue Min_played $0,100]\subset\mathbb{R}$ Total minutes played at each interval; proxy for physical wear.
Disciplinary Risk Card_Y $\{0,1\}\subset\mathbb{Z}$ Indicates if player received a yellow card; higher risk increases substitution priority.
Form Trend Momentum $1,1]\subset\mathbb{R}$ Short-term rate of change in performance (3-interval moving average). Negative values denote declining form.
Contextual Fatigue (Age) Age $15,45]\subset\mathbb{N}$ Player’s age on match day; modulates fatigue effects.
Tactical Role Position CategoricalCategorical Role label; used for role-aware normalization.
Offensive Contribution (Goals) Goals $\mathbb{N}$ Cumulative goals scored; high values indicate offensive importance.
Offensive Contribution (Assists) Assists $\mathbb{N}$ Cumulative assists; complements goal contribution.
@@
5.4.2 Spearman Correlation Matrix
@@
6 Fuzzy Control System
@@
6.1 Definition of Fuzzy Variables
@@ 6.2 Membership Functions
6.3 Rule Base
The rule base was streamlined to remove noise and focus on Intensifying or attenuating the player’s performance using other contextual events as variables. The system avoids oscillation by ignoring minor fluctuations and focuses on critical states. The logic integrates position and stats directly. :
@@ 6.3 Rule Base
ID Rule Logic
R01 – Untouchable Star IF (P_cum is High/VeryHigh) THEN (Modifier = Negative).
R02a – Critical Fatigue IF (P_cum Low/VeryLow) AND (Min High) THEN (Modifier = Large Positive).
R02b – Early Fatigue IF (P_cum Low/VeryLow) AND (Min Medium) THEN (Modifier = Medium Positive).
R03 – Defensive Risk IF (Defender AND YellowCard Yes) THEN (Modifier = Medium Positive).
R04 – Rapid Decline IF (P_cum Low/VeryLow AND Momentum Falling) THEN (Modifier = Large Positive).
R07 – Positive Momentum IF (P_cum High/VeryHigh AND Momentum Rising) THEN (Modifier = Medium Negative).
R08 – Ineffective Forward IF (Forward AND P_cum Low AND Goals None) THEN (Modifier = Large Positive).
R09 – Striker Under Pressure (1) IF (Forward AND P_cum Low/Med AND Momentum Falling) THEN (Modifier = VeryLarge Positive).
(2) IF (Forward AND P_cum Low/Med AND Min High) THEN (Modifier = VeryLarge Positive).
R10 – Invisible Playmaker IF (Midfielder AND P_cum Low AND Assists None) THEN (Modifier = Med–Large Positive).
R11 – Creator Bonus IF (Assists Some OR Goals Some) THEN (Modifier = Medium Negative).
IF (Assists Many OR Goals Many) THEN (Modifier = Large Negative).
R12 – Veteran Fatigue IF (Age Veteran AND Min High) THEN (Modifier = Medium Positive).
R13 – Young Talent Protection IF (Age Young AND (Goals Some/Many OR Assists Some/Many)) THEN (Modifier = VeryLarge Negative).
R14 – Goal Protection IF (Goals Some) THEN (Modifier = Medium Negative).
IF (Goals Many) THEN (Modifier = Large Negative).
R15 – Neutral State IF (P_cum Medium AND Momentum Stable) THEN (Modifier = Zero).
6.4 Inference and Final Calculation
@@
$Baseline=100\times(1.0-P_{cum})$ (1)
$P_{final}=\text{clip}\left(Baseline+(Modifier\times\alpha),\ 0,\ 100\right)$ (2)
@@ 7.1 Adherence to Technical Commission Decisions
Player Analysis Slice($t$) $P_{final}$ in the slice Rank at $t$ Predominant Fuzzy Logic Real Decision
@@
Fagner 45’ – 85’ 100.0 1st Fatigue Not Substituted
@@ 9 Conclusion
However, limitations remain. The current reliance on event data as a proxy for physical fatigue (minutesPlayed) serves as an estimation rather than a direct measurement. Future iterations of this framework will prioritize the integration of biometric telemetry and GPS tracking data to refine the fatigue inputs. Furthermore, we envision the development of a Neuro-Fuzzy (ANFIS) architecture, allowing the rule base to be dynamically calibrated through feedback from professional analysts, effectively creating a hybrid system that learns from expert consensus while maintaining algorithmic objectivity.

### v3

@@ Abstract
In elite soccer, substitution decisions entail significant financial and sporting consequences yet remain heavily reliant on intuition or predictive models that merely mimic historical biases. This paper introduces a Fuzzy Logic-based Decision Support System (DSS) designed for real-time, prescriptive game management. Unlike traditional Machine Learning approaches that encounter a “predictive ceiling” by attempting to replicate human behavior, our system audits performance through an objective, rule-based inference engine. We propose a methodological advancement by reformulating the PlayeRank metric into a Cumulative Mean with Role-Aware Normalization, eliminating the play-time exposure bias inherent in cumulative sum models to enable accurate intra-match comparison. The system integrates this refined metric with physiological proxies (fatigue) and contextual variables (disciplinary risk modulated by tactical role) to calculate a dynamic Substitution Priority ($P_{final}$). Validation via a case study of the 2018 FIFA World Cup match between Brazil and Belgium demonstrates the system’s ecological validity: it not only aligned with expert consensus on executed substitutions (e.g., Gabriel Jesus) but, crucially, identified high-risk scenarios ignored by human decision-makers. Specifically, the model flagged the “Fagner Paradox”, a maximum-priority defensive risk, minutes before a critical yellow card, and detected the “Lukaku Paradox,” where an isolated assist masked a severe drop in participation. These results confirm that Fuzzy Logic offers a transparent, explainable, and superior alternative to “black-box” models for optimizing real-time tactical decisions.
@@ 2 Literature Review
Recent work supports this perspective. SchmidtLilloBustos2024 demonstrate that player influence fluctuates meaningfully over a match due to tactical reconfiguration, opponent pressure, and physical factors, advocating for temporal windowing to capture such dynamics. Following this insight, the present work also adopts 5-minute temporal slices, but advances the methodology by adding a role-aware percentile cumulative mean across slices. This modification preserves interpretability while enabling direct comparability between players with different playing times, eliminating residual exposure bias and allowing the performance curve to reflect true temporal variability.
A second methodological challenge concerns the interpretation of imprecise or uncertain information in decision-making contexts. The literature highlights fuzzy logic as an effective tool for this purpose. For instance, Huarachi-Macuri et al. LACCEI2023 demonstrate that fuzzy systems are suitable for transforming quantitative descriptors into interpretable linguistic assessments (e.g., Stamina, Agility). Within tactical modeling, Marliere2017 uses fuzzy control systems to arbitrate between tactical states under uncertainty. These contributions motivate the role of fuzzy logic in the present system: integrating performance indicators and contextual factors into a Substitution Priority score in a manner compatible with the inherent uncertainty of in-game decision making.
@@
3 Methodology
3.1 Description of the Problem
3.1.1 Problem Formulation
Tactical decision-making in elite soccer operates under high uncertainty and significant financial stakes GE_PremierLeague; ESPN_Mundial. Despite the proliferation of granular data, the specific process of substitution decisions remains predominantly intuitive or reliant on descriptive statistics that fail to capture real-time performance decay. This creates two distinct problems: the Metric Exposure Bias and the Predictive Ceiling.
The Metric Exposure Bias: Existing frameworks for player evaluation, such as the original PlayeRank Pappalardo2019, typically utilize cumulative sum formulations. While effective for post-match rankings, these metrics introduce a temporal bias during live games: a player’s score monotonically increases with playing time, regardless of their minute-by-minute efficiency. This mathematical structure masks declining performance, as a fatigued player performing poorly in the $80^{th}$ minute may still have a higher total score than a high-impact substitute, rendering standard metrics insufficient for real-time substitution decisions.
The Predictive Ceiling of Machine Learning: Current computational approaches to substitutions rely heavily on Supervised Machine Learning (SML) models trained on historical data. As noted by Mohandas et al. mohandas, these models achieve a prediction accuracy plateau of approximately 70%. This ”ceiling” exists because SML models are designed to mimic human behavior, thereby learning and replicating the cognitive biases of coaches (e.g., status quo bias or delaying defensive changes). Consequently, these models validate historical decisions rather than optimizing future outcomes, failing to identify necessary substitutions that deviate from conservative human norms.
Therefore, the core problem this study addresses is the lack of an objective, prescriptive auditing tool that can quantify intra-match performance decay without exposure bias and signal tactical risks independently of historical human tendencies.
3.1.2 Research Hypotheses
To address the formulated problem, this study tests the following hypotheses:
  • H1 (Metric Hypothesis): Reformulating performance evaluation from a cumulative sum to a Cumulative Mean with Role-Aware Normalization effectively eliminates exposure bias, allowing the system to detect negative momentum and performance drops that are mathematically invisible in additive models.
  • H2 (Prescriptive Hypothesis): A Fuzzy Logic-based Decision Support System, acting as an objective auditor of performance variables (Performance, Fatigue, Discipline), can identify high-risk tactical scenarios (e.g., ”invisible” defensive liabilities) earlier and more accurately than human intuition or predictive models trained on historical behavior.
  • H3 (Contextual Hypothesis): The modulation of disciplinary risk by tactical role (e.g., weighting a yellow card more heavily for a defender than a forward) significantly alters the substitution priority score, aligning the algorithm’s output with expert tactical consensus in critical defensive scenarios.
3.2 Dataset
@@
3.3 Data Integration and Pre-processing
@@
3.3.1 Stage 1: Temporal Performance Metric Generation
@@
Subsequently, to ensure metric interpretability independent of absolute match intensity, the raw score is normalized against the historical distribution of the agent’s specific tactical role. This transformation converts the unbounded scalar playerank_fatia_raw into the Role-Aware Percentile (playerank_fatia_percentil), ensuring that a ”High” score represents the same statistical rarity for a Defender as it does for a Forward.
@@
$\mathrm{playerank\_acumulativo\_media\_percentil}\ \ \textbf{=}\ \ \frac{1}{t}\sum_{i=1}^{t}\mathrm{playerank\_fatia\_percentil}^{(i)}$
@@
3.3.2 Stage 2: Contextual and Demographic Enrichment
@@
  • momentum_rate: a measure of short-term performance trend, calculated as the difference between the current($t$) and the previous(t1t-1) 5-minute playerank_acumulativo_media_percentil slices.
@@
3.3.3 Stage 3: Final Cleaning and Role-Aware Normalization
@@
  • removal of out-of-field periods: Periods where it was identified that the player was substituted, or entered as a substitute, were removed to ensure that the metric remained accurate, not applied to goalkeepers.
@@ 3.3.4 Data Validation
3.4 Exploratory Data Analysis
@@ 3.4.1 Descriptive Overview
(playerank...percentil) 0.501 symmetrical distribution
@@ 3.4.2 Univariate Distributions
3.4.3 Role-Based Event Distributions
To further interpret positional behavior, the dataset was aggregated by player position to estimate empirical probabilities of key in-game events (goals, assists, and yellow cards).
The empirical distributions align with established tactical expectations:
@@ 3.4.3 Role-Based Event Distributions
3.4.4 Variable Definition, Independence and Correlations
Variable Definition and Selection
@@ 3.4.4 Variable Definition, Independence and Correlations
Dimension Variable (ID) Mathematical Set Description and Relevance Technical Performance P_cum $0,1]\subset\mathbb{R}$ Cumulative performance percentile per match and position; main indicator of global efficiency. Fatigue Min_played $0,100]\subset\mathbb{R}$ Total minutes played at each interval; proxy for physical wear. Disciplinary Risk Card_Y $\{0,1\}\subset\mathbb{Z}$ Indicates if player received a yellow card; higher risk increases substitution priority. Form Trend Momentum $1,1]\subset\mathbb{R}$ Short-term rate of change in performance (3-interval moving average). Negative values denote declining form. Contextual Fatigue (Age) Age $15,45]\subset\mathbb{N}$ Player’s age on match day; modulates fatigue effects. Tactical Role Position $Categorical$ Role label; used for role-aware normalization. Offensive Contribution (Goals) Goals $\mathbb{N}$ Cumulative goals scored; high values indicate offensive importance. Offensive Contribution (Assists) Assists $\mathbb{N}$ Cumulative assists; complements goal contribution.
@@
Spearman Correlation Matrix
@@
3.5 Fuzzy Control System
@@
3.5.1 Definition of Fuzzy Variables
@@ 3.5.2 Membership Functions
3.5.3 Rule Base
The rule base was streamlined to remove noise and focus on Intensifying or attenuating the player’s performance using other contextual events as variables. The system avoids oscillation by ignoring minor fluctuations and focuses on critical states. The logic integrates position and stats directly.
@@ 3.5.3 Rule Base
ID Rule Logic R01 – Untouchable Star IF (P_cum is High/VeryHigh) THEN (Modifier = Negative). R02a – Critical Fatigue IF (P_cum Low/VeryLow) AND (Min High) THEN (Modifier = Large Positive). R02b – Early Fatigue IF (P_cum Low/VeryLow) AND (Min Medium) THEN (Modifier = Medium Positive). R03 – Defensive Risk IF (Defender AND YellowCard Yes) THEN (Modifier = Medium Positive). R04 – Rapid Decline IF (P_cum Low/VeryLow AND Momentum Falling) THEN (Modifier = Large Positive). R07 – Positive Momentum IF (P_cum High/VeryHigh AND Momentum Rising) THEN (Modifier = Medium Negative). R08 – Ineffective Forward IF (Forward AND P_cum Low AND Goals None) THEN (Modifier = Large Positive). R09 – Striker Under Pressure (1) IF (Forward AND P_cum Low/Med AND Momentum Falling) THEN (Modifier = VeryLarge Positive). (2) IF (Forward AND P_cum Low/Med AND Min High) THEN (Modifier = VeryLarge Positive). R10 – Invisible Playmaker IF (Midfielder AND P_cum Low AND Assists None) THEN (Modifier = Med–Large Positive). R11 – Creator Bonus IF (Assists Some OR Goals Some) THEN (Modifier = Medium Negative). IF (Assists Many OR Goals Many) THEN (Modifier = Large Negative). R12 – Veteran Fatigue IF (Age Veteran AND Min High) THEN (Modifier = Medium Positive). R13 – Young Talent Protection IF (Age Young AND (Goals Some/Many OR Assists Some/Many)) THEN (Modifier = VeryLarge Negative). R14 – Goal Protection IF (Goals Some) THEN (Modifier = Medium Negative). IF (Goals Many) THEN (Modifier = Large Negative). R15 – Neutral State IF (P_cum Medium AND Momentum Stable) THEN (Modifier = Zero).
3.5.4 Inference and Final Calculation
@@
$\text{Baseline}=100\times(1.0-P_{\text{cum}})$ (1)
$P_{\text{final}}=\text{clip}\left(\text{Baseline}+(\text{Modifier}\times\alpha),\ 0,\ 100\right)$ (2)
@@ 4.1 Adherence to Technical Commission Decisions
Player Analysis Slice($t$) $P_{\text{final}}$ in the slice Rank at $t$ Predominant Fuzzy Logic Real Decision
@@
Fagner 45’-85’ 100.0 1st Fatigue Not Substituted
@@ 5.3 Case Study: The ”Fagner Paradox” as Proof of Utility
6 Contributions
This study advances the field of sports analytics and decision support systems through the following novel contributions:
  • Methodological Innovation: The ”Cumulative Mean” Metric(H1).
    We propose a mathematical reformulation of the PlayeRank framework Pappalardo2019. Unlike the original cumulative sum model, which inherently introduces an exposure bias favoring players with longer field time, we developed a Cumulative Mean with Role-Aware Normalization. This formulation eliminates residual bias, enabling accurate intra-match comparison (PlayeRank up to a specific moment in the match) and the detection of real-time performance decay, a capability absent in previous static metrics.
  • Domain Adaptation: From Robotics to Human Management(H2).
    A fundamental distinction of this work lies in its application scope. While previous studies like Marliere2017 employed fuzzy logic to dictate immediate navigation actions, mapping geometric inputs (e.g., distance, angle) directly to robotic actuator commands, this study transposes the logic to the tactical-managerial level. We utilize intrinsic player variables to model a temporal risk score. Unlike the reactive approach common in robotics, our model integrates the match’s historical evolution (via Cumulative Mean and Momentum), allowing the system output to capture performance decay over time rather than just the instant state.
  • Hybrid ”Contextual Tuner” Architecture(H2).
    We introduce a novel inference architecture where the Fuzzy Control System (FCS) functions as a contextual modifier applied to a robust statistical baseline, rather than a standalone decision-maker. By weighting the fuzzy output ($\alpha=0.25$), the system balances objective performance data with situational urgency (e.g., disciplinary risk or fatigue), ensuring stability and preventing the oscillation often found in pure rule-based systems.
  • Codification of Tactical Heuristics(H3).
    The system formalizes subjective coaching intuitions into computable variables, specifically through the introduction of a ”Momentum” metric (rate of performance change) and Role-Modulated Risk (weighting disciplinary sanctions differently for defenders vs. forwards). This allows the model to identify subtle ”invisible” risks, such as the ”Fagner Paradox” (high defensive liability without statistical failure), which purely statistical models fail to capture.
  • Paradigm Shift: Prescriptive Auditing over Prediction(H2).
    We demonstrate that traditional Supervised Machine Learning models encounter a ”predictive ceiling” ($\approx 70\%$) by training to mimic human biases (e.g., status quo bias, sunk cost fallacy). Our approach shifts to a prescriptive audit, prioritizing objective utility over historical patterns. This was empirically validated by identifying high-priority substitution risks minutes before they negatively impacted the match result, outperforming human decision-making in critical scenarios.
@@ 7 Conclusion
7.1 Limitations and Future Works
While the proposed Fuzzy Control System (FCS) demonstrates significant utility as a prescriptive auditing tool, certain limitations inherent to the current design and data sources must be acknowledged.
First, the reliance on minutes_played as a proxy for physical fatigue assumes a linear accumulation of load, serving as an estimation rather than a direct physiological measurement. This approach fails to account for inter-individual physiological differences (e.g., genetics), variations in game intensity, such as the volume of high-speed running, or both, like the metabolic energy expended by individual players. Consequently, the model may underestimate the fatigue of a player subjected to high defensive pressure or overestimate the wear of a player in a low-block system.
Second, the static nature of the fuzzy rule base relies on heuristically defined expert knowledge. While this ensures explainability, it may not perfectly generalize across divergent tactical philosophies (e.g., a high-pressing team versus a possession-based team) without manual recalibration of the membership functions or rule weights ($\alpha$). Furthermore, the ecological validation, although robust within the high-stakes context of the 2018 World Cup, requires expansion to a longitudinal analysis to quantify the system’s long-term consistency.
Future iterations of this framework will address these gaps through three primary advancements:
  1. 1.
    Integration of Biometric Telemetry: Prioritizing the ingestion of GPS, accelerometer data and bioindicators to refine fatigue inputs, shifting from temporal proxies to metrics based on internal and external load.
  2. 2.
    Neuro-Fuzzy Adaptation (ANFIS): We envision the development of an Adaptive Neuro-Fuzzy Inference System (ANFIS). This architecture would allow the rule base and membership functions to be dynamically calibrated through supervised learning on historical expert decisions, effectively creating a hybrid system that learns from elite consensus while maintaining algorithmic objectivity.
  3. 3.
    Tactical Profile Matching: Expanding the system’s output to suggest not only who should be substituted, but also identifying the optimal replacement profile from the bench based on the specific tactical gap detected (e.g., replacing a fatigued “runner” with a fresh “playmaker”).